OpenStudy (anonymous):

Given an arithmetic series with a1 = 70 and the sum of the first 12 terms = 576. What is d?

OpenStudy (mimi_x3):

Arithmetic series formula: \[T_n = a+(n-1)d\] Given: a_ 1 = 70 ; and the first 12 terms is 576 so... \[ T_{12} = 70 + (12-1)d = 576\] Are you able to do it now? :)

OpenStudy (anonymous):

should i add 72+12/576

OpenStudy (mimi_x3):

You are solving for \(d\) so you expand \((12-1)d\) And where did \(72\) come from? :/

OpenStudy (anonymous):

I meant 70

OpenStudy (mimi_x3):

Well .. \( 70+(12-1)d = 576\) Can you expand the \((12-1)d\) first?

OpenStudy (anonymous):

11

OpenStudy (anonymous):

how do i expand it